Circular No. 7914

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002de IN NGC 6104 M. Papenkova and W. D. Li, University of California at Berkeley, re-

port the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova (mag ∼ 16.0) on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on June 1.4 and 2.4 UT. SN 2002de is located at α = 16^{h}16^{m}30^{s}.38, δ = +35^{o}42′30′′.2 (equinox 2000.0), which is 4′′.0 west and 1′′.0 north of the nucleus of NGC 6104. KAIT images taken on May 13.4 (limiting mag ∼ 19.0) and 23.4 (limiting mag ∼ 18.0) showed nothing at this position.

COMET C/2002 C1 (IKEYA-ZHANG) G. Cremonese, Istituto Nazionale di Astrofisica (INAF) and Osserva-

torio Astronomico, Padova; A. Boattini, M. T. Capria, M. C. De Sanctis, and G. D’Abramo, Istituto di Astrofisica Spaziale e Fisica Cosmica, CNR, Rome; and A. Buzzoni, INAF and Telescopio Nazionale Galileo (TNG), report observing the sodium distribution in the coma of comet C/2002 C1 on Apr. 20.17 UT, using the 3.5-m TNG reflector (+ high-resolution spectrograph SARG; resolving power 43000; spatial coverage 26′′) with a narrow-band filter to isolate the sodium-D lines. Two long-slit spectra (one parallel and the other perpendicular to the sun-comet vector) show cometary sodium emissions clearly visible due to the comet’s geocentric velocity of −8.3 km/s. In the first spectrum, the sodium emissions are clearly asymmetric with respect to the slit center, corresponding to the nuclear region, extending to the edge of the slit in the tail direction (∼ 4000 km). Even in the second spectrum, a slight asymmetry is visible in the sodium emissions with stronger intensity toward the southwest, suggesting the presence of a sodium jet that is most likely related to a dust feature.

V838 MONOCEROTIS D. P. K. Banerjee and N. M. Ashok, Physical Research Laboratory

(PRL), Ahmedabad, report that near-infrared spectroscopy of V838 Mon at 1.08–2.35 μm, obtained with the Mt. Abu 1.2-m telescope (+ PRL Near-Infrared Imaging Spectrometer + HgCdTe NICMOS3) on May 2 and 14, show several special and unusual emission lines seen in the K band, which match well (in wavelength and strength) the expected lines from TiI at 2.2639, 2.245, 2.2280, 2.2239, 2.2011, 2.19034, and 2.1789 μm. The Na I doublet at 2.2056 and 2.2084 μm is also seen in emission. The Ti I lines were are also found in data from Apr. 9, but they are strongest on May 2.

2002 June 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green